OpenStudy (anonymous):

log3(X-5)+log3(x+3)=2

OpenStudy (anonymous):

well according to log rules, when u have a + of logs with the same base you can multiply the terms so log3[(x-5)(x+3)]=2 and then convert it to e so (x-5)(x+3)=3^2 solve for x :)

OpenStudy (anonymous):

so LN base e both sides?

OpenStudy (anonymous):

Nevermind, I figured it out...thanks!!!

OpenStudy (anonymous):

the answer is 6 :)